OpenStudy (anonymous):

Solve √(3x+2) =√(2x+3)

OpenStudy (anonymous):

Lol.

OpenStudy (anonymous):

If I help you, don't say you don't understand and disagree with my answer. Lol.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Omg. This is easy. x = 1. Lol

OpenStudy (anonymous):

3(1) + 2 = 5 2(1) + 3 = 5

OpenStudy (anonymous):

sqrt(5) = sqrt(5)

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

ty